OpenStudy (anonymous):

∫sin(-2x-1)cos(2x-1)

OpenStudy (zzr0ck3r):

sin(2x) = 2sin(x)cos(x) sin(2x)/2 = sin(x)cos(x) sin(2(-2x-1))/2 = sin(-2x-1)(cos(-2x-1) sin(-4x-2) = sin(-2x-1)cos(-2x-1) integrate the left hand side

OpenStudy (mandre):

Slight mistake there zzr0ck3r

OpenStudy (zzr0ck3r):

fix it:)

OpenStudy (mandre):

I don't know integration. I'm just saying you put cos(-2x-1) and it's cos(2x-1). I'm learning integration next week lol.

OpenStudy (zzr0ck3r):

ahh I forgot the 2 sin(-4x-2)/2 = sin(-2x-1)cos(-2x-1)

OpenStudy (zzr0ck3r):

ohh woops, good call. well that changes everything....lol

OpenStudy (psymon):

Wonder if the extra negative inside of the sin angle is supposed tobethere. *kicks site* finally working again.

OpenStudy (zzr0ck3r):

yeah im thinking the same thing, or if the one on the right is missing it

OpenStudy (zzr0ck3r):

seems to odd...

OpenStudy (mandre):

I was thinking the same thing Psymon.

OpenStudy (mandre):

lol

OpenStudy (mandre):

Can't you use the identity for sin(a)cos(b)?

OpenStudy (psymon):

No reason to.

OpenStudy (zzr0ck3r):

there are lots of ways to do it, I was going for the easy one.

OpenStudy (psymon):

I thought the (1/2)sin2x one reeked of win, so let's stick to that.

OpenStudy (zzr0ck3r):

then we are done \[\int sin(-2x-1)cos(-2x-1)dx=\frac{1}{2}\int sin(-4x-4)dx = \frac{1}{8}cos(-4x-2)\]

OpenStudy (zzr0ck3r):

u=2x+1?

OpenStudy (zzr0ck3r):

not that it matters, im just trying to learn it:)

OpenStudy (mandre):

slight mistake there lol

OpenStudy (dumbcow):

oh yeah the signs dont match

OpenStudy (mandre):

You with us now dumbcow lol

OpenStudy (zzr0ck3r):

we are thinking his question has a typo, and he medaled me so maybe so...no idea

OpenStudy (psymon):

Which is why we just do some science and throw in some magic extra negative and pretend they are the same angle.

OpenStudy (dumbcow):

yeah http://www.wolframalpha.com/input/?i=integrate+-sin%282x%2B1%29cos%282x-1%29+dx prob need to expand whole thing using sum/difference identity

OpenStudy (zzr0ck3r):

that's odd...sin(2) lol

OpenStudy (psymon):

Haha, I like the random sin(2)

OpenStudy (mandre):

Alarm bells ringing lol

OpenStudy (dumbcow):

if this is from a regular textbook....all bets on typo :)

OpenStudy (psymon):

Or a cruel person giving out hw?

OpenStudy (zzr0ck3r):

@Aoi please put us out of our misery

OpenStudy (psymon):

青いさん D:

OpenStudy (mandre):

Same to you psymon

OpenStudy (psymon):

x_x Okay, leaving then

OpenStudy (zzr0ck3r):

lol

OpenStudy (mandre):

Note to self: learn pinyin.

OpenStudy (mandre):

'Aoi' sounds like something a samurai would say before comitting hara kiri.

OpenStudy (anonymous):

Take u=cos(2x-1)

OpenStudy (anonymous):

du=-2sin(2x-1)

OpenStudy (anonymous):

U get sin(-(2x-1))=-sin(2x-1) Multiple Numerator and denominator by 2 and take (1/2) outside!!

OpenStudy (zzr0ck3r):

we are in no rush just hit enter when you are done:)

OpenStudy (anonymous):

\[{1/2}\int\limits u du\]

OpenStudy (anonymous):

Integrate the function........ Substitute u=cos(2x-1) and u get the answer,I guess!!

OpenStudy (anonymous):

I hope thats right @zzr0ck3r @Mandre @Psymon @Aoi